OpenStudy (anonymous):

Find the inverse of the function: f(x)=1+e^x/1-e^x

OpenStudy (anonymous):

\[\frac{ 1+e ^{x} }{ 1-e ^{x} }\] I know the opposite of e is ln, but thats about it

OpenStudy (anonymous):

is it\[f(x)=\frac{1+e^x}{1-e^x}\]

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

\[x=\frac{1+e^y}{1-e^y}\] find e^y

OpenStudy (anonymous):

here we swapped x with y

OpenStudy (anonymous):

I get that, but do I just multiply by ln on the top and bottom?

OpenStudy (anonymous):

before lets multiply both sides by 1-e^y

OpenStudy (anonymous):

oh to get it off the bottom?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do i get:

OpenStudy (anonymous):

do i get:\[x(1-e ^{x})=(1+e ^{x})(1-e ^{x})\]

OpenStudy (anonymous):

i mean those exponents to be y's

OpenStudy (anonymous):

right hand side does not have the 1-e^x part\[\frac{ 1+e^x }{ \cancel{1-e^x }}\cancel{1-e^x}\]

OpenStudy (anonymous):

oh gotcha!

OpenStudy (anonymous):

now i mutiply by ln on both sides?

OpenStudy (anonymous):

so we have\[x(1-e^y)=1-e^y\]

OpenStudy (anonymous):

open bracket then take take e^y to the same side

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

I got as far as your last equation and then I multiplied by ln on both sides and get confused on what can cancel

OpenStudy (anonymous):

do you get this\[x(1-e^y)=1-e^y\]\[x-xe^y=1-e^y\]

OpenStudy (anonymous):

hey we never multiplied by ln both sides

OpenStudy (anonymous):

oh gotcha, distribute the x through, ok

OpenStudy (anonymous):

so can you take xe^y to the right and 1 to the left

OpenStudy (anonymous):

alrighty

OpenStudy (anonymous):

\[x+1=-xe ^{y}-e ^{y}\]

OpenStudy (anonymous):

notice that its now +xe^y not -xe^y change of sign

OpenStudy (anonymous):

oh ok, mistake on just copying it to here

OpenStudy (anonymous):

ok lol so \[x-1=e^y(x-1)\] amazingly canceling x-1 both sides we have \[1=e^y\]

OpenStudy (anonymous):

how can we just cancel both sides?

OpenStudy (anonymous):

oh by dividing, nevermind

OpenStudy (anonymous):

yes so\[\ln e^y=\ln 1\]

OpenStudy (anonymous):

so y=ln1 , which is y=0. thanks!

OpenStudy (anonymous):

wait somthing went wrong

OpenStudy (anonymous):

Its no biggie. I think i get the basic idea now. Thanks though

OpenStudy (anonymous):

i can always just ask my teacher for sure on monday

OpenStudy (anonymous):

I think its because its actually - on top and + on the bottom of the main funciton. No sense doing it over though

OpenStudy (anonymous):

I miscopied it from the book i guess

OpenStudy (anonymous):

\[x=\frac{ 1+e^y }{ 1-e^y }\] \[x-e^y=1\color{red}{+}e^y\] we wrote - earlier so its + \[x-1=xe^y+e^y\] \[1-x=e^y(1+x)\] \[e^y=\frac{ 1-x }{ 1+x}\]

OpenStudy (anonymous):

yep! that makes the books answer now if i just ln both sides. thanks so much!

OpenStudy (anonymous):

\[y=\ln \left(\frac{ x-1 }{ x+1 }\right)\]

OpenStudy (anonymous):

thanks so much. i fanned you and gave you best response

OpenStudy (anonymous):

yw